Calculate the [OH−] of aqueous solution with the pH = 11.0.
Express your answer using two significant figures.


[OH−] = ???
1
Expert's answer
2015-11-17T08:58:17-0500
The sum of pH and pOH is equal 14, from it we can find pOH.
pH + pOH = 14
pOH = 14 - pH = 14-11 = 3
From the formula for pOH calculation we can estimate the concentration of [OH-].
pOH = - log [OH-]
[OH-] = 1x10^-3 M = 0.001 M

Answer:
[OH-] = 0.001 M